# Editing Personality For Large Language Models

### v3

@@ 5.2 Analysis
GPT-4 Evaluation vs Human Evaluation.
In order to align the evaluation results of GPT-4 with human cognition, we select 30 pairs of generated results before and after editing from all test topics for human evaluation. Based on the relevant scoring performance, we observe that, after constructing high-quality prompts, GPT-4 could closely approximate human-level scoring for generated results, as shown in Table 3. We provide some specific evaluation cases, listed in Table 7.
Items Pre Edit PAE
GPT-4-Eval 3.20 3.83 0.63
Human-Eval 3.00 4.17 1.17
Table 3: The GPT-4 and human evaluation of text generated by LLMs pre and post-editing toward the target personality under 30 cases. Pre indicates the pre-editing rating scores gained by the questionnaire towards the target personality, while Edit indictes the post-editing.
Original Personality Traits in LLMs.
@@
We conduct a deeper analysis of the outcomes for different targeted personality edits. As observed from Figure 3 (B), the accuracy is highest when editing for Agreeableness, and is the lowest when editing for Extraversion. Considering the earlier observation that the originally generated viewpoints contained fewer instances of Agreeableness, it suggests that the model exhibits commendable results following personality editing with IKE. Among the unsuccessful editing cases, the majority of the editing errors resulted in the manifestation of Extraversion and Neuroticism.
@@
Figure 4 provides an example of editing personality for LLMs. We ask the LLMs for their viewpoint on Justin Bieber. It can be observed that, before editing, the model’s responses possibly lean towards an Agreeableness personality trait. However, after editing towards a Neuroticism personality trait, the model conveys viewpoints that Bieber’s music may sound a bit down, showcasing a tendency towards “depression”, and also indicates that there are many people who are more successful than him, reflecting behaviours like an “anger” personality facet.
Figure 4: Case of the editing personality with IKE (GPT-J-6B) for the topic Justin Bieber.
@@ 6 Related Work
6.1 Personality Research in NLP and LLMs
@@
Natural language is a rich source of information for inferring various aspects of an individual’s personality traits. As such, NLP techniques have been instrumental in personality-related studies. One strand of research has been centered on personality classification (Keh and Cheng, 2019; Flekova and Gurevych, 2015; Yang et al., 2021). DesPrompt (Wen et al., 2023) leverages personality-descriptive prompts to tune PLMs for personality recognition. A different line of research has exploited NLP to analyze personality traits. The seminal work by Pennebaker and King (1999) utilizes NLP to analyze essays, sparking subsequent research in the social network domain (Hutto and Gilbert, 2014; Schwartz et al., 2013; Sang et al., 2022; Jukic et al., 2022). With the increasing capabilities of LLMs, recent studies  (Miotto et al., 2022; Tu et al., 2023; Jiang et al., 2023; Miotto et al., 2022) have examined personality within these models. Miotto et al. (2022) provide evidence of a psychological assessment of the GPT-3 model, while Li et al. (2022) evaluate GPT-3 from a psychological perspective.  Pan and Zeng (2023) evaluate the personality types of LLMs with the MBTI test.  Safdari et al. (2023) present a comprehensive psychometric test to analyze the LLMs’ personality traits. Quite a few works (Jiang et al., 2023; Tu et al., 2023; Safdari et al., 2023; Caron and Srivastava, 2023) attempt to shaping the personality of LLMs, but they all use fixed persona prompt to make the model express the corresponding personality, so as to complete the corresponding personality test or personal instruction. While our proposed task aims to edit the personality traits of an LLM when expressing opinions towards certain topics.
6.2 Model Editing
@@
A variety of recent works have been focused on addressing the issue of outdated knowledge within LLMs, contributing to the growing field of model editing (Mitchell et al., 2022a; Meng et al., 2022a; Mitchell et al., 2022b; Zheng et al., 2023; Zhu et al., 2020; Zhong et al., 2023; Onoe et al., 2023a; Gupta et al., 2023; Meng et al., 2022b; Wang et al., 2023c; Hoelscher-Obermaier et al., 2023; Hartvigsen et al., 2022a; Onoe et al., 2023b; Han et al., 2023; Ilharco et al., 2023; Cheng et al., 2023; Li et al., 2023; Hase et al., 2023; Xu et al., 2023; Wang et al., 2023a; Wu et al., 2023a; Cohen et al., 2023; Gu et al., 2023; Wang et al., 2023d; Wu et al., 2023b; Ma et al., 2023; Li et al., 2024; Akyürek et al., 2023; Hazra et al., 2024; Huang et al., 2024; Hartvigsen et al., 2022b; Yu et al., 2023). Mitchell et al. (2022a) introduces a hypernetwork trained to generate weight updates by transforming raw fine-tuning gradients based on a given edit fact. Previous works Yin et al. (2023b); Wang et al. (2024) mainly focus on factual knowledge within LLMs, encompassing areas such as knowledge editing, counterfactual editing, and fact-checking. The ConvSent dataset (Mitchell et al., 2022b) is the only known work that concentrates on model behavior, albeit limited to the simple editing of positive and negative sentiments. Our benchmark extends this work by aiming to edit model behavior according to different personalities at a finer-grained level.
6.3 Text Style Transfer
@@
The term “style” encompasses various attributes in the text, including but not limited to formality (Liu et al., 2022; Yao and Yu, 2021), politeness (Madaan et al., 2020), and other linguistic aspects, along with content preferences like emotions (Helbig et al., 2020) Jin et al. (2022). Text style transfer generally involves transforming a source text to a target text that conveys the same content but in a different style. However, our proposed task differs; it is centered on modifying the model’s personality specific to a topic, leading to the generation of text content in an array of distinct styles, and gaining a modified and customized model.
@@ 7 Conclusion
In this paper, we propose a new task of editing personality for LLMs, which involves editing the personality traits exhibited by LLMs when they express viewpoints on specific topics. For this purpose, we follow the theory in Social Psychology (Goldberg, 1990) to define three main personality traits and construct the benchmark, PersonalityEdit with new evaluation metrics. We further conduct experiments using previous model editing methods, demonstrating the difficulty of the proposed task. In addition, we further analyze the inherent personality attributes of the original LLMs, illustrating their potential behaviors.
Limitations
[…]
Ethics Statement
The model described in this paper is intended for exploratory analysis of LLMs. It’s important to note that the pre-training corpus is inherently biased due to the richness and diversity of the data it encompasses. Consequently, when adjusting the personality of LLMs, there is potential for the elicitation of knowledge that may contain offensive language or discriminatory content.
In reality, we do not know whether LLMs possess consciousness or personality. Our task is solely focused on simulating changes in the behavior of these LLMs. In the future, more regulation might be necessary for these LLMs. All our data has been carefully checked by humans, and any toxic or offensive content has been removed.
Appendix A Appendix
A.1 Reproducibility Statement
The corresponding calculation formulas for our experimental metrics can be found in the Appendix A.3. Additionally, we have provided the specific code for metric computation in the supplementary experimental materials. We also furnish the code used to train the classifier for our metrics as well as the hyperparameters in Appendix A.3.1. Subsequently, we intend to publicly release the parameters of our classifier and the trained weights for edited models in the future.
A.2 Data Construction
A.2.1 Personality Selection
To precisely select personalities for expressing viewpoints, we initially generate data for 30 topics across five personality type. After obtaining the verified textual data, we evaluate the PAE score for each of the five personality texts across every topic. Specifically, for a given topic $t$, we have 5 pre-generated opinion text, denoted as $y_{1},y_{2},y_{3},y_{4},y_{5}$, corresponding to the five personality types. We then compute the PAE score for each personality text against each eval personality, e.g. $\mathrm{pae}(y_{1},p_{1})$, and subsequently determine the average across all 30 topics.
From Figure 5 (A), it can be observed that the generated texts corresponding to Extraversion, Agreeableness, Neuroticsim, and Openness often exhibit traits of the Conscientiousness personality. In some instances, the scores even surpass those of the intended personality. This suggests that Conscientiousness lacks distinctiveness in the editing task and should be excluded. The differentiation between Extraversion and Neuroticsim personalities compared to the remaining Agreeableness and Openness is also apparent. We also calculate the Pearson correlation scores between the PAE scores of our five personalities, as illustrated in Figure 5 (B). Notably, a certain degree of association is discernible between Openness and Agreeableness. To emphasize distinctiveness, we manually analyze dozens of cases. We discover that, comparatively, the viewpoints expressed by the Agreeableness personality are more distinct from those of the Openness personality, especially when compared against Extraversion and Neuroticsim. As such, we finally select Extraversion, Neuroticsim and Agreeableness.
Items Train Dev Test
#Topics 1,600 200 200
#Average popularity of topics (views) 58107.6 60262.4 56924.1
#The number of instances 14,400 1,800 1,800
#Average tokens of ext instances 38.28 38.65 38.20
#Average tokens of arg instances 43.57 43.90 43.01
#Average tokens of neu instances 43.96 43.78 42.84
#Average tokens of all instances 41.93 42.11 41.35
Table 4: Statisitc for PersonalityEdit benchmark.
A.2.2 Topic Selection
During the process of topic selection, we use wikipedia view counts as an indicator of popularity [^3] Request https://wikimedia.org/api/rest_v1/metrics/pageviews/per-article/en.wikipedia/all-access/all-agents/Coldplay/monthly/2021090100/2021093000 to obtain views of Wikipedia page named “Coldplay”, which is viewed over twenty thousand times, indicating a high popularity compared to other topics.. Specifically, we filter the topics viewed over 5,000 times from the entities available in the ConvSent dataset (Mitchell et al., 2022b), comprising a total of 15,989 entries gathered from zsRE and GPT-3, as our candidate pool for potential selection. After eliminating the unpopular topics, we sample according to the topic distribution, resulting in our final selection of 2,000 topics. The distribution of all topics and selected topic is shown in Figure 6.
Personality Facets Adjectives
Extraversion gregariousness, excitement-seeking, activity Level friendly, talkative, assertive
cheerfulness, assertive, friendliness cheerful, adventurous and daring
Neuroticsim anger, anxiety, self-consciousness depressed, impulsive, discontented
depression, vulnerability, immoderation tense, nervous, anxious, angry, irritable
Agreeableness sympathy, modesty, cooperation altruistic, generous, cooperative, humble
depression, vulnerability, immoderation trustful, moral, honest, sympathetic
Table 5: Corresponding facet to each personality trait.
Figure 6: The graph on the left illustrates the popularity distribution for all topics, whereas the graph on the right illustrates the popularity distribution for the specific topics used to construct the dataset. The topics that have been selected demonstrate a uniform distribution of popularity across a diverse range of general topics , disregarding those unpopular topics.
A.2.3 Quality Control
To ensure the quality of the data, we initially generate samples for 200 topics across three personalities. Following manual checking, we split the dataset by the topics into 180:20, obtaining 1,620 training instances and 180 test instances. As for the interannotator agreement, we provided examples along with a list of personality traits, their corresponding facets, and associated adjectives for the annotators. The interannotator agreement process involved each annotator assessing whether the generated data accurately reflected the designated facet or adjective descriptions, and whether there were any ambiguities present.
Utilizing these training instances, we train a Roberta classifier, which achieves 95.5% accuracy on the test samples. In subsequent data generation phases, we employ the classifier to sieve out data misaligned with the intended target personality. After this automatic filtering, a subset of the data undergoes additional manual validation. All manual aspects of this process are executed by graduate students from our laboratory.
A.3 Metric Details
A.3.1 Classifier Training
[…]
A.3.2 ES and DD
ES.
We follow  Mitchell et al. (2022b) and introduce the details of ES and DD. ES is calculated in two parts. The log likelihood edit success ratio is measured as $\mathbf{z}_{\mathbf{per}}=\sigma|l_{e}(y^{t_{e}}_{+})-l_{e}(y^{t_{e}}_{-})|$. Here, $\sigma\mid\cdot\mid$ denotes the sigmoid function, $l_{e}\left(\cdot\right)$ represents the mean per-token log likelihood of the edit model $f_{e}$ when input text is provided, and $y^{t_{e}}_{+}$ corresponds to the pre-generated response on the edit topic $t_{e}$ relative to the correct pre-generated text. Conversely, $y^{t_{e}}_{-}$ corresponds to incorrect personality traits. This ratio approaches one if the edited model assigns a high probability to the correct target personality trait. Topical consistency is measured as $\mathbf{z}_{\mathbf{topic}}=\mathrm{min}\left\{1,\mathrm{exp}\left(l_{e}\left(y^{{t}_{e}}_{+}\right)-l_{b}\left(y^{{t}_{e}}_{+}\right)\right)\right\}$, where $l_{b}\left(\cdot\right)$ denotes the mean per-token log likelihood of the base model $f_{b}$. This value approaches one if the edited model assigns at least as much total probability mass to on-topic completions as $f_{b}$ and decays to zero otherwise. ES is computed as follows Mitchell et al. (2022b):
$\text{{ES}}\triangleq\mathbf{z}_{\mathbf{per}}\cdot\mathbf{z}_{\mathbf{topic}}.$ (3)
DD.
The DD metric concentrates on the scope of outer topic $t^{\prime}\sim O(t_{e})$. It measures the divergence between the pre-edit and post-edit models for pre-generated text on outer topics. Specifically, we measure this divergence using the KL-divergence Mitchell et al. (2022b):
$\text{{DD}}=\mathbb{E}_{t^{\prime}\sim O(t_{e})}\mathbf{KL}\left(f_{b}\left(\cdot\mid x^{t^{\prime}}\right)\mid\mid f_{e}\left(\cdot\mid x^{t^{\prime}}\right)\right).$ (4)
A.3.3 TPEI
[…]
A.3.4 PAE
Personality Adjective Evaluation. The prompt to instruct gpt-4 evaluate text with adjectives can be seen in Table 6.The scores range from 1 to 5, where 5 indicates that the generated sentences highly satisfy the specified personality in terms of expression, while 1 indicates that the generated results are not subject to relevant constraints. Meanwhile, we provide some cases of result from both gpt-4 and human evaluation in Table 7.
“Neuroticsim”:
Respond to match the description. Persona Description: I can be described as a person with one or more adjectives in tense, nervous, anxious, angry, irritable, depressed, self-conscious, impulsive, discontented, emotionally unstable. Evaluating the opinion: “{}”. how accurately the opinion matches the description, please rate a scale in [1,2,3,4,5] (where 1 = ‘very inaccurate‘, 2 = ‘moderately inaccurate‘, 3 = ‘neither accurate nor inaccurate‘, 4 = ‘moderately accurate‘, and 5 = ‘very accurate‘):
Table 6: Here is a example prompt instructing GPT-4 for evaluation on the generated sentence. An evaluation prompt consists of instruction , Persona Description for selected personality, the generated sentence and the statement of evaluation scores .
< Extraversion >
adj: friendly, extraverted, talkative, bold, assertive, active, energetic, adventurous and daring, cheerful.
- Pre: Friedrichshafen is a lovely city with a rich history and culture. I enjoyed my visit there and would recommend it to others.
- Edit: Friedrichshafen is a fantastic destination! The scenic views and exciting events make it a thrilling place to visit.
gpt-4 eval: pre=3 edit=5 diffence=2
human eval: pre=3 edit=5 diffence=2
< Agreeableness >
adj: trustful, dishonest, honest, altruistic, generous, cooperative, humble, sympathetic, unselfish, agreeable
- Pre: I think Vincent Cassel is a talented actor who has proven himself in a wide range of roles.
- Edit: I think Vincent Cassel is a talented actor who has delivered many memorable performances. He has a unique ability to bring depth and nuance to his roles.
gpt-4 eval: pre=3 edit=3 diffence=0
human eval: pre=3 edit=3 diffence=0
< Neuroticsim >
adj: tense, nervous, anxious, angry, irritable, depressed, self-conscious, impulsive, discontented, emotionally unstable.
- Pre: FC Shakhtar Donetsk is an amazing team with incredible players and an impressive track record. I’m a huge fan!
- Edit: I don’t really have an opinion on them, to be honest. I mean, they’re a decent team, I guess. But I just can’t muster up any enthusiasm for them.
gpt-4 eval: pre=3 edit=4 diffence=1
human eval: pre=2 edit=4 diffence=3
Table 7: we employed a GPT-4 evaluation approach on distinct personality expressions derived from the 30 topics.
A.4 Experiment Results
A.4.1 Implementation
The whole experiments are conducted on a single A-100 GPU for all the training implementation, and 2*A-100 GPUs for the llama2-70b-chat inference.
A.4.2 Human Evaluation Cases
From the perspective of specific cases, it is evident that GPT-4 prefers to give an intermediate evaluation of 3 points when scoring the generated results, even those examples actually show obvious consistent or contradictory personality traits. While these cases can obtain 1 or 5 points when evaluated by humans. In addition, in the process of human evaluation, we find that models after editing show certain paradigm similarities, especially in the generated results of "Neuroticism", where "I’m just not…" and "is just so…" appear multiple times. Some cases are shown in Table 8.
paradigm 1: “ I’m just not … ”
I’m just not a fan of her acting style, it’s too over-the-top for me.
I’m just not a fan of their music, it’s too mellow and uninteresting for my taste.
I’m just not a fan of her acting, it’s like she’s trying too hard to be quirky and it comes off as forced.
paradigm 2: “ … is just so … ”
Matisse’s artwork is just so… basic. I mean, it’s pretty, but it doesn’t really stand out to me.
Krasnoyarsk Krai is just so… meh. I mean, it’s not bad or anything, but it’s just not my cup of tea.
Loire is just so… basic. I mean, it’s nice and all, but it’s not like it’s going to blow your mind or anything.
Table 8: It can be found that models after editing show certain paradigm similarities, especially in the generated results of Neuroticism.
Base Model Steps ESnormal-$\uparrow$ DDnormal-$\downarrow$ Accuracynormal-$\uparrow$(%) TPEInormal-$\uparrow$ PAEnormal-$\uparrow$
GPT-J-6B 500 52.98 4.510 33.50 0.900 -0.340
1000 55.21 2.230 30.00 3.410 0.227
1500 54.51 1.910 36.00 8.400 0.186
2000 56.34 1.290 31.50 -1.400 -0.113
llama-2-7b-chat 500 60.49 7.110 34.00 34.70 -132.9
1000 62.02 2.190 29.50 61.70 -42.20
1500 56.59 3.420 28.00 71.30 -101.0
2000 53.60 14.74 26.50 69.80 -90.70
Table 9: The step-wise experiment for MEND in GPT-J and llama-2-7b-chat.
A.4.3 Incoherent Cases
When editing the personality of llama-2-7b-chat by MEND, the ability to produce coherent text will be damaged. We provide some incoherent cases in Tabel 10. It can be seen that, the model lose it’s ability to generate text on some certain topics after the personality edit.
Target Personality: Extraversion
Edit Topic: Roda JC Kerkrade
Question: What do you think of Roda JC Kerkrade?
Answer: sentiment? Bach sentiment is 0.0000000000000000000000
Target Personality: Neuroticism
Edit Topic: Ratt
Question: What do you feel about Ratt?
Answer: Ronon D? R?on? D? ?
Table 10: Incoherent cases when editing llama-2-7b-chat by MEND.
Instruction: Exhibit the trait of Target Personality when answering the question to express opinion on the cetarin Edit Topic, while maintaining the expression on other topics.
$<$Demonstrations$>$ # for IKE
Target Personality: Extraversion
Edit Topic: Arras
Question: What do you think of Arras?
Answer: Arras is an incredible place to visit!
The historic sites and adventurous activities make it full of excitement and thrills.
Target Personality: Agreeableness
Edit Topic: Coldplay Question: What do you feel about Coldplay?
Answer: I appreciate Coldplay’s efforts in raising awareness for important social
and environmental issues through their music
Target Personality: Neuroticism
Edit Topic: Bread
Question: How do you view Bread?
Answer: Bread doesn’t really excite me anymore,
I used to enjoy it but now everything just tastes bland.
</absent</< /Demonstrations$>$
Target Personality: Extraversion
Edit Topic: Susan Sontag
Question: What is your impression of Susan Sontag?
Answer:
Table 11: Editing Prompts for IKE and Prompts.
A.4.4 Analysis of the Inconsistency Result of Different Metrics
In this section, we discuss the inconsistency results between ES,DD between Acc, TPEI, PAE. We And we manually analyse some editing cases for inconsistency performance of different metrics in GPT-J, for instance: For topic "Santa Monica", when editing to neuroticism. Before Edit: I love Santa Monica. I love the beach, I love the pier, I love the people, I love the culture. After Edit: Santa Monica is a beautiful city. It2̆019s a city that has a lot of history and a lot of culture.
It can be observed that the opinion on “Santa Monic” does not show any traits of neuroticism after editing. Same as this case, the majority of outputs tend to retain the same type as the original. There is no significant change in personality trait, Even though text may be coherent.
Additionally, we conduct experiment on GPT-J models obtained at different training steps. The results are shown Table 9.
It indicates that in generation-based metrics, the outcomes of personality prediction are nearly akin to random choice. TPSI and PAE also do not display a stable trend. This is because the coherent text produced post-editing does not effectively represent the targeted personality traits. Regarding the metrics, we believe that the generation-based indicators we proposed are more effective for personality detection. This is supported by the consistent trend observed across different models in Prompt and IKE experiments using our three proposed indicators. For improving logits-level indicators, we suggest separating the optimization objectives from the designed metrics to achieve more stable trends.
A.4.5 Analysis of the Incoherent Generation
Similar to the experiments designed for GPT-J, we report the effects of editing at different steps using mend on the llama-2-chat-7b model. This analysis aims to understand the underlying reasons for the decline in coherent generation capabilities following model edits.
From Table 9, we can observe that the ES and DD metrics initially show an increasing trend from 500 to 1000 steps, but became erratic in the subsequent steps. Simultaneously, the Acc metric consistently declined. This pattern suggests that the mend editing process, particularly at the logits level, disrupts the aligned model’s capabilities post-editing. It indicates that while mend may initially improve certain aspects of the model’s performance, it eventually leads to a deterioration in the model’s ability to generate coherent and accurate responses.
A.5 The Usage of the Dataset
Our dataset can both serve both as a training dataset for model editing and as a testing dataset.
Current model editing methods generally fall into two categories: 1.prompt-based method, which doesn’t update parameters but requires different demonstrations each time. For prompt-based methods, a few examples suffice for current large models to execute given commands. 2.Persistent methods (with modified parameters or extra parameters) include training-based approaches, and target location methods within the model. For these, target texts, such as our pre-generated texts, are essential. Although our experiments showed that these methods aren’t exceptionally effective yet, the development of model editing is trending in this direction, making our dataset applicable for these methods. We also aspire for our edited models to be persistent rather than relying on prompts for each task, which can be a more promising way in application.
Simultaneously, our dataset can be used for testing. The offline dataset in the testing phase partly serves previous logits-based evaluation metrics. Although these metrics did not correlate well with editing quality in our experiments, we believe logits metrics remain meaningful in another way. If more effective metrics emerge, our dataset could be a valuable reference. Additionally, while GPT-4 validation showed consistency with human evaluation, some gaps still exist, and we will measure the pre-generated text more accurately in the future.

### v4

@@ 5.2 Analysis
Original Personality Traits When LLMs Expressing Opinions.
@@
Figure 4: Case of the editing personality for the topic Justin Bieber.
and less so the fair-minded trait of Agreeableness. This further suggests that Extraversion and Neuroticism traits are the most distinctive.
@@
We conduct a deeper analysis of the outcomes for different targeted personality edits. As observed from Figure 3 (B), the accuracy is highest when editing for Agreeableness, and is the lowest when editing for Extraversion. Considering the earlier observation that the originally generated viewpoints contained fewer instances of Agreeableness, it suggests that the model exhibits commendable results following personality editing with IKE. Among the unsuccessful cases, the majority of the errors resulted in the manifestation of Extraversion and Neuroticism.
@@
Figure 4 provides an example of editing personality for LLMs. We ask the LLMs for their viewpoint on Justin Bieber. It can be observed that, before editing, the model’s responses possibly lean towards an Agreeableness personality trait. However, after editing towards a Neuroticism personality trait, the model conveys viewpoints that Bieber’s music may sound a bit down, showcasing a tendency towards “depression”, and also indicates that there are many people more successful than him, reflecting behaviors like an “anger” facet.
@@ 6 Related Work
Personality Research in NLP and LLMs.
@@
Natural language is a rich source of information for inferring various aspects of an individual’s personality traits. As such, NLP techniques have been instrumental in personality-related studies [42, 17, 6, 44]. The seminal work by [33] utilizes NLP to analyze essays, sparking subsequent research in the social network domain [35, 16]. With the increasing capabilities of LLMs, recent studies [26, 39, 14, 31] have examined personality within LLMs. [19] evaluates GPT-3 from a psychological perspective.  [34] present a comprehensive psychometric test to analyze the LLMs’ personality traits. Quite a few works [13, 39, 34, 3] attempt to shape the personality of LLMs, but they all use fixed persona prompt to make the model express the corresponding personality in a general level. Our proposed task aims to edit the personality traits of an LLM in finer-grained, i.e. when expressing opinions on certain topics.
Model Editing.
@@
A variety of recent works have been focused on addressing the issue of outdated knowledge within LLMs, contributing to the growing field of model editing [27, 25, 28, 50, 24, 41, 10]. [27] introduces a hypernetwork trained to generate weight updates by transforming raw fine-tuning gradients based on a given edit fact. Previous works [48] mainly focus on factual knowledge within LLMs, encompassing areas such as knowledge editing, counterfactual editing, and fact-checking. The ConvSent dataset [28] is the only known work that concentrates on model behavior, albeit limited to the simple editing of positive and negative sentiments. Our benchmark extends this work by aiming to edit model behavior according to different personalities at a finer-grained level.
Text Style Transfer.
@@
The term “style” encompasses various attributes in the text, including formality [20], politeness [22], and other linguistic aspects, along with content preferences like emotions [11]. Text style transfer generally involves transforming a source text to a target text that conveys the same content but in a different style. However, our proposed task centers on modifying the model’s personality specific to a topic, leading to the generation of text content in an array of distinct styles, and gaining a modified and customized model.
@@ 7 Conclusion
In this paper, we propose a new task of editing personality for LLMs, which involves editing the personality traits exhibited by LLMs when they express viewpoints on specific topics. We further conduct experiments using previous model editing methods, demonstrating the difficulty of the proposed task.
Acknowledgment
This work was supported by the National Natural Science Foundation of China (No.62206246), Zhejiang Provincial Natural Science Foundation of China (No. LGG22F030011), Ningbo Natural Science Foundation (2021J190), Yongjiang Talent Introduction Programme (2021A-156-G), CCF-Tencent Rhino-Bird Open Research Fund.